# Social learning moderates the tradeoffs between efficiency, stability, and equity in group foraging

### v1

@@ Acknowledgements
Research was sponsored by the DEVCOM Analysis Center and was accomplished under Contract Number W911QX-23-D0002. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the DEVCOM Analysis Center or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@
After testing the case of 10 agents searching for the targets with the same distribution as that in Fig. 2, we can see that the optimal $\rho$ still exists, and it corresponds to the maximal time ratios of exploitation. However, the shape of efficiency curves are flatter compared to cases with 50 agents in the main text. This occurs because the smaller number of agents limits the maximum instantaneous efficiency that can be achieved by social learning and the following targeted walk, thereby diminishing the advantages of collective search. As a result, the efficiency curves the efficiency curves show less variation and are less dependent on $\rho$. This suggests that the model is more effective for larger groups. The group size constraint on instantaneous efficiency also leads agents to use a larger $\rho$ to attract more peers to achieve higher instantaneous efficiency, which, in turn, increases the optimal $\rho$. Additional videos are available at https://www.youtube.com/playlist?list=PLgRFM9nAjJRwoZvCGBAdCIE-BYNgPmSuV.

### v2

@@ Acknowledgements
Research was sponsored by the DEVCOM Analysis Center and was accomplished under Contract Number W911QX-23-D0002. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the DEVCOM Analysis Center or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.
@@
We thank Janson W. Burton, Lingfei Wu, and Hirokazu Shirado for helpful discussion.
@@
This Supplementary Information provides additional analyses and robustness checks that support the main text. Section S1 examines how the target-clustering parameter $\beta$ affects search efficiency and behavioral time allocations. Section S2 reports time-resolved (instantaneous) search efficiency to clarify how target depletion influences the efficiency curves and the optimal social learning range $\rho$. Section S3 evaluates the effect of group size on the existence and location of the optimal $\rho$. Finally, Section S4 details the construction of the proximity-based social network induced by social learning and defines the network metrics used to quantify connectivity and the emergence of collective structure over time.
@@
After testing the case of 10 agents searching for the targets with the same distribution as that in Fig. 2, we can see that the optimal $\rho$ still exists, and it corresponds to the maximal time ratios of exploitation. However, the shape of efficiency curves are flatter compared to cases with 50 agents in the main text. This occurs because the smaller number of agents limits the maximum instantaneous efficiency that can be achieved by social learning and the following targeted walk, thereby diminishing the advantages of collective search. As a result, the efficiency curves the efficiency curves show less variation and are less dependent on $\rho$. This suggests that the model is more effective for larger groups. The group size constraint on instantaneous efficiency also leads agents to use a larger $\rho$ to attract more peers to achieve higher instantaneous efficiency, which, in turn, increases the optimal $\rho$.
@@
S4 Network Metrics
To quantify how the social learning range $\rho$ shapes interaction structure, we define an implicit, time-dependent proximity network among agents. At each timestep $t$, each agent is treated as a node, and an undirected edge is placed between agents $i$ and $j$ if their Euclidean separation satisfies
$\|\mathbf{x}_{i}(t)-\mathbf{x}_{j}(t)\|<\rho.$ (S.1)
This produces a sequence of graphs $G(t)$ over the course of each realization. We compute standard network statistics from $G(t)$ to characterize how information-sharing range maps onto effective connectivity, local transitivity, and fragmentation.
The degree of agent $i$ at time $t$ is $k_{i}(t)$, the number of neighbors within distance $\rho$. Unless otherwise stated, the degree time series shown in Supplementary Fig. 6A is the median (over realizations) of the network-averaged degree at each timestep.
Local transitivity is quantified using the (local) clustering coefficient $C_{i}(t)$, defined for nodes with $k_{i}(t)\geq 2$ as
$C_{i}(t)=\frac{2T_{i}(t)}{k_{i}(t)\left(k_{i}(t)-1\right)},$ (S.2)
Supplementary Figure 6: Network dynamics across social learning ranges. Median (over independent realizations) time series of proximity-network metrics for a range of social learning radii $\rho$. At each timestep, an edge is defined between two agents if their separation is smaller than $\rho$. (A) Median node degree, showing the growth of effective local connectivity as interactions accumulate and as $\rho$ increases. Median clustering coefficient, quantifying the emergence of locally transitive interaction neighborhoods. (C) Median number of connected components $N_{C}$, measuring network fragmentation and the transition from many disconnected groups to a single connected network as $\rho$ increases. (D) Median largest component size $S_{\mathrm{max}}$, capturing the growth of the giant component and the onset of near-global connectivity at large $\rho$. Together, the panels illustrate how increasing $\rho$ accelerates the collapse of fragmentation and the formation of a highly connected, locally clustered interaction network over time.
where $T_{i}(t)$ is the number of triangles that include node $i$. The network clustering coefficient $C(t)$ is taken as the mean of $C_{i}(t)$ over all nodes with $k_{i}(t)\geq 2$ (nodes with $k_{i}<2$ contribute zero by convention). The time series in Supplementary Fig. 6B shows the median (over realizations) of $C(t)$. Values near 0 indicate weakly transitive neighborhoods, whereas values near $1$ indicate highly transitive local structure.
A connected component is defined as a maximal set of nodes in which every pair is connected by at least one path. Let $N_{C}(t)$ be the number of connected components at time $t$. We also define the component sizes $\{s_{1}(t),s_{2}(t),\dots,s_{N_{C}(t)}(t)\}$, where each $s_{m}(t)$ is the number of nodes in component $m$. The largest component size is
$S_{\max}(t)=\max_{m}s_{m}(t).$ (S.3)
Supplementary Fig. 6C reports the median (over realizations) of $N_{C}(t)$, and Supplementary Fig. 6D reports the median (over realizations) of $S_{\max}(t)$. Together, these quantities capture the transition from a fragmented network (large $N_{C}$, small $S_{\max}$) at small $\rho$ to near-global connectivity (small $N_{C}$, large $S_{\max}\approx N_{A}$) at large $\rho$.
In addition to time-resolved metrics, we compute case-level component summaries used in Fig.3 EF of the main text. Specifically, for each realization we compute (i) the average component size $S$ (obtained by averaging component sizes over time), and (ii) the average number of components $N_{C}$. We then aggregate across realizations to form the PDFs shown in Fig.3 EF and to compute the case-averaged trends $\langle S\rangle$ and $\langle N_{C}\rangle$ versus $\rho$. These component-based statistics provide a compact network-level interpretation of $\rho$: increasing $\rho$ merges components, increases typical connected set sizes, and accelerates the emergence of a giant component, thereby increasing effective information transmission.
For each $\rho$, we run 2000 independent realizations. For time-resolved plots (Supplementary Fig. 6), we compute the relevant network statistic at each timestep in each realization and then report the median across realizations at each time to reduce sensitivity to outliers. This yields robust trajectories of how connectivity and fragmentation develop during the search process.